OpenStudy (anonymous):

I really need help with this. Its frustrating me. Earnest opens a savings account with $5,000. He deposits $900 each year into the account that compounds quarterly and has a 0.75% interest rate. What will his account total be in 5 years? $9,771.94 $12,649.12 $10,706.30 $23,515.14

OpenStudy (anonymous):

Please please help

OpenStudy (anonymous):

c

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

im on google right now i copy pasted your question

OpenStudy (anonymous):

can you paste how that person did that problem?

OpenStudy (tkhunny):

Once you have comprehensible information, you can just piece it together. "account that compounds quarterly and has a 0.75% interest rate" What does that mean? Is the Annual rate 0.75%? Is the Quarterly rate 0.75%? If it had said, "compounds at a quarterly interest rate of 0.75%" that would have been clear. Let's assume it's a quarterly rate of 0.75% i = 0.0075 -- The quarterly interest rate. r = 1+i = 1.0075 -- The Quarterly accumulation factor Now building! (5000r^4 + 900) - One year ((5000r^4 + 900)r^4 + 900) - Two years (((5000r^4 + 900)r^4 + 900)r^4 + 900) - Three years ((((5000r^4 + 900)r^4 + 900)r^4 + 900)r^4 + 900) - Four years (((((5000r^4 + 900)r^4 + 900)r^4 + 900)r^4 + 900)r^4 + 900) - Five years Not very pretty, is it? A little algebra... 5000r^20 + 900r^16 + 900r^12 + 900r^8 + 900r^4 + 900 - Five years - Much nicer. 4100r^20 + (900r^20 + 900r^16 + 900r^12 + 900r^8 + 900r^4 + 900) - Five years - Interesting. 4100r^20 + 900(r^20 + r^16 + r^12 + r^8 + r^4 + 1) - Five years - Better. This is the fun one! \(4100r^{20} + 900\left(\dfrac{1 - r^{24}}{1 - r^{4}}\right)\) - Five years - THAT was awesome. You DO need to know how to create the sum of a finite geometric series. Note: This is building your own formula for the specific circumstance. It is not required to memorize what formula to use. Get good at creating them. Note: Not everyone agrees with my opinion on this matter. Your algebra has to be super sharp and that seems like to great a requirement for some.

OpenStudy (anonymous):

where did the 4100 come from? and dont you have to reduce the interest rate 0.0075/4

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

"where did the 4100 come from?" The initial payment is 5000. Since all the other payments were 900, I grabbed 900 of the initial deposit for the big, fancy formula, leaving 4100 by itself at the beginning. " and dont you have to reduce the interest rate 0.0075/4" You must pay better attention. I addressed this issued quite directly up front. Your problem statement is not clear on this issue. I ASSUMED that 0.75% was the quarterly rate. If it is somehow clear that this is not the case, you will have to deal with that. The formulation is the same. This is why I used 'r' and 'i', so that you could substitute any applicable value.

OpenStudy (anonymous):

ok. thanks.

OpenStudy (tkhunny):

You can see it clearly on the second to last equation. One of the 900s and the 4100 have the same exponent on their respective accumulation factors - indicating they occur at the same time.

OpenStudy (anonymous):

can you help with another one?

OpenStudy (anonymous):

What was the answer?

OpenStudy (anonymous):

@chelseachels